OpenStudy (anonymous):

Solve exactly. \[\log_{3} \](2x-3) = 0

OpenStudy (anonymous):

\[Solve exactly. \log_{3} \](2x-3) = 0\]

OpenStudy (anonymous):

2x-3 = 3^0 2x-3 = 1 x = 2

OpenStudy (anonymous):

the first line comes from the definition of logarithm